The Parker Company sold a $1,000 par value bond that has 20 years to maturity and a 7.00% annual coupon rate with coupons paid semiannually. The bond currently sells for $950, and the company’s tax rate is 33%. What is the AFTER TAX component cost of debt for use in the WACC calculation?

Enter the answer as a decimal with four places of precision (i.e. 0.1234).
Answer #1

Coupon = (0.07 * 1000) / 2 = 35

Number of periods = 20 * 2 = 40

Yield to maturity = 7.48608%

Keys to use in a financial calculator:

2nd I/Y 2

FV 1000

PV 950

N 40

PMT 35

CPT I/Y

After tax cost of debt = 0.0748608 (1 - 0.33)

After tax cost of debt = 0.0502